### v1

@@ A.1 Out-of-domain Evaluation
In Table 9, we report additional out-of-domain results, and we evaluate the models on the CC3M dataset, which includes image-caption pairs sourced from web repositories. The results show consistency with those observed on the nocaps and VizWiz datasets, reported in the main paper. Notably, employing PAC-S++ as reward consistently enhances semantic richness while preserving fluidity during generation, as demonstrated by the higher CIDEr scores than the one achieved by the model optimized via CLIP-S reward. This improvement is evident across both ViT-B/32 and ViT-L/14 backbones, further confirming the effectiveness of our training strategy and its generalization capabilities to out-of-domain datasets.
@@
Backbone Reward C CLIP-S PAC-S++ RefCLIP-S RefPAC-S++
- 22.8 0.643 0.653 0.638 0.688
CIDEr 27.9 0.655 0.663 0.657 0.698
CLIP-S 0.6 0.710 0.691 0.579 0.675
PAC-S++ 9.5 0.702 0.697 0.621 0.686
Ref-CLIP-S 21.1 0.679 0.678 0.660 0.702
ViT-B/32 RefPAC-S++ 24.6 0.676 0.686 0.661 0.709
- 27.1 0.653 0.665 0.648 0.699
CIDEr 34.8 0.665 0.672 0.676 0.713
CLIP-S 0.8 0.726 0.708 0.609 0.690
PAC-S++ 13.8 0.715 0.708 0.639 0.698
Ref-CLIP-S 27.2 0.696 0.693 0.678 0.719
ViT-L/14 RefPAC-S++ 32.0 0.688 0.697 0.681 0.724
A.2 Additional Qualitative Results

### v2

@@ A.1 Out-of-domain Evaluation
In Table 11, we report additional out-of-domain results, and we evaluate the models on the CC3M dataset, which includes image-caption pairs sourced from web repositories. The results show consistency with those observed on the nocaps and VizWiz datasets, reported in the main paper. Notably, employing PAC-S++ as reward consistently enhances semantic richness while preserving fluidity during generation, as demonstrated by the higher CIDEr scores than those achieved by the model optimized via CLIP-S reward. This improvement is evident across both ViT-B/32 and ViT-L/14 backbones, further confirming the effectiveness of our training strategy and its generalization capabilities to out-of-domain datasets.
@@
Backbone Reward C CLIP-S PAC-S++ RefCLIP-S Polos RefPAC-S++
- 22.8 0.643 0.653 0.638 0.265 0.688
CIDEr 27.9 0.655 0.663 0.657 0.253 0.698
CLIP-S 0.6 0.710 0.691 0.579 0.300 0.675
PAC-S++ 9.5 0.702 0.697 0.621 0.318 0.686
Ref-CLIP-S 21.1 0.679 0.678 0.660 0.303 0.702
ViT-B/32 RefPAC-S++ 24.6 0.676 0.686 0.661 0.301 0.709
- 27.1 0.653 0.665 0.648 0.304 0.699
CIDEr 34.8 0.665 0.672 0.676 0.273 0.713
CLIP-S 0.8 0.726 0.708 0.609 0.342 0.690
PAC-S++ 13.8 0.715 0.708 0.639 0.359 0.698
Ref-CLIP-S 27.2 0.696 0.693 0.678 0.342 0.719
ViT-L/14 RefPAC-S++ 32.0 0.688 0.697 0.681 0.343 0.724
Figure 7: Sample failure cases of a captioning model fine-tuned with different metrics as reward.
A.2 Limitations
While the proposed metric demonstrates strong performance across several benchmarks, it is important to acknowledge its limitations and examine failure cases that suggest directions for future improvement. Notably, our model inherits some constraints from the underlying CLIP architecture, as both the textual and visual encoders in our framework are based on CLIP backbones. Although we apply task-specific finetuning through a positive augmentation strategy, the core architectural limitations remain.
On the textual side, the main limitation stems from the fixed token length (i.e. 77 tokens), which can lead to truncation and loss of semantic information in longer textual descriptions. This might restricts the model ability to fully leverage context-rich or multi-sentence inputs. On the visual side, the use of a fixed input resolution can cause the loss of fine-grained details, especially in high-resolution or complex scenes. Moreover, since CLIP generates a global image embedding, the model may struggle with tasks requiring fine-grained recognition or explicit modeling of spatial relationships between objects.
Failure Cases. To provide some insights for future research, we also include representative failure cases in Fig. 7, when employing PAC-S++ as reward to fine-tune a captioning model. While our approach already shows a clear reduction in repetitive captions, as shown in Table 9, some challenges such as hallucination and object miscounting still persist. These are known limitations across captioning systems and are difficult to fully eliminate – even with strong reward signals like PAC-S++. In particular, our model can misjudge spatial relationships in the absence of explicit depth or perspective information, such as describing a plane as “between” two buildings when it is actually flying behind them. It can also struggle with accurately counting objects, especially when they are occluded or only partially visible. Subjective attributes, like distinguishing between “tall” and “short” grass, often remain ambiguous even to human observers, further complicating the task. Additionally, the model may occasionally hallucinate plausible yet nonexistent background details. Although these errors are relatively rare and often subtle, they highlight ongoing challenges in the field and point toward meaningful directions for future improvement.
A.3 Additional Qualitative Results